Assume total productivity is given as Q = 1600L + +50L2 - 1.5L3

Determine the marginal productivity of the 10th worker and the average productivity of the first 10 workers.

Also, show the work for determining the number of labor to maximize productivity. How many units of our goods will be produced?
Answer #1

We have Q = 1600L + +50L2 - 1.5L3

This gives MPL = dQ/dL = 1600 + 100L - 4.5L^2 and APL = Q/L = 1600 + 50L - 1.5L^2

MPL (L = 10) = 1600 + 100*(10) - 4.5*(10^2) = 2150 units

APL (L = 10) = 1600 + 50*10 - 1.5*(10^2) = 1950 units

To maximize productivity, MPL should be 0

This gives 1600 + 100L - 4.5L^2 = 0.

This happens at L = 33 where MPL becomes 0

At L = 33, a total of 1600*33 + 50*33^2 - 1.5*(33^3) = 53344.50 units are produced.